### v3

# In-Context Learning with Iterative Demonstration Selection

Abstract
Spurred by advancements in scale, large language models (LLMs) have demonstrated strong few-shot learning ability via in-context learning (ICL). However, the performance of ICL has been shown to be highly sensitive to the selection of few-shot demonstrations. Selecting the most suitable examples as context remains an ongoing challenge and an open problem. Existing literature has highlighted the importance of selecting examples that are diverse or semantically similar to the test sample while ignoring the fact that the optimal selection dimension, i.e., diversity or similarity, is task-specific. Based on how the test sample is answered, we propose Iterative Demonstration Selection (IDS) to leverage the merits of both dimensions. Using zero-shot chain-of-thought reasoning (Zero-shot-CoT), IDS iteratively selects examples that are diverse but still strongly correlated with the test sample as ICL demonstrations. Specifically, IDS applies Zero-shot-CoT to the test sample before demonstration selection. The output reasoning path is then used to choose demonstrations that are prepended to the test sample for inference. The generated answer is followed by its corresponding reasoning path for extracting a new set of demonstrations in the next iteration. After several iterations, IDS adopts majority voting to obtain the final result. Through extensive experiments on tasks including reasoning, question answering, and topic classification, we demonstrate that IDS can consistently outperform existing ICL demonstration selection methods.
In-Context Learning with Iterative Demonstration Selection
Chengwei Qinthanks: Correspondence to Chengwei Qin <chengwei003@e.ntu.edu.sg> and Aston Zhang <az@astonzhang.com>, Aston Zhang, Chen Chen, Anirudh Dagar,Wenming Ye Nanyang Technological University,Amazon Web Services
1 Introduction
Figure 1: Illustration of in-context learning (ICL) on topic classification. A frozen large language model directly generates the topic ‘Technology’ for the test sample ‘OpenAI …’ by taking the demonstrations and the test sample as input.
With the recent advancements in scaling up model parameters, large language models (LLMs) showcase promising results on a variety of few-shot tasks through in-context learning (ICL), where the model is expected to directly generate the output of the test sample without updating parameters. This is achieved by conditioning on a manually designed prompt consisting of an optional task description and a few demonstration examples (Brown et al., 2020). Fig. 1 shows an example describing how LLMs perform ICL on the topic classification task. Given a few text-topic pairs as demonstrations, ICL combines them with the test sample as input, to the LLM for inference. The output, i.e., ‘Technology’, is generated by the model autoregressively without any parameter updates.
Despite the effectiveness, the performance of ICL has been shown to be highly sensitive to the selection of demonstration examples (Zhao et al., 2021). Different sets of demonstrations can yield performance ranging from nearly random to comparable with state-of-the-art models (Gao et al., 2021; Lu et al., 2022). To alleviate the above issue, researchers in ICL have proposed a number of methods to select a set of examples as few-shot demonstrations (Rubin et al., 2022; Liu et al., 2022; Li and Qiu, 2023; Wang et al., 2023b; Li et al., 2023a; Ma et al., 2023; An et al., 2023b). However, for LLMs for which parameters or detailed output distributions are not available (Sun et al., 2022), it is still a common practice to randomly select examples or select examples that are semantically similar to the test sample as demonstrations, i.e., considering diversity or similarity. While several approaches investigate the combination of similarity and diversity when prompting with explanations, exploring compositional generalization, or choosing examples for annotation (Ye et al., 2023b; An et al., 2023a; Su et al., 2023), it is not yet clear how to determine and leverage the optimal dimension for different tasks in ICL and how the rationale for answering the query benefits the balance between these two dimensions.
Actually, the optimal dimension for selecting demonstration examples is task-specific. As we will show in Section 4, the diversity dimension is superior to the similarity dimension on CommonsenseQA while the similarity dimension outperforms the diversity dimension on AGNews and BoolQ. Thus, it is unreasonable to claim that one dimension is consistently better than the other across different tasks. To fully leverage the merits of both dimensions, we propose Iterative Demonstration Selection (IDS) for ICL (Fig. 2) by utilizing how the test sample is answered. IDS can iteratively select demonstration examples that are diverse but still have a strong correlation with the test sample through zero-shot chain-of-thought reasoning (Zero-shot-CoT) (Kojima et al., 2022). Specifically, Zero-shot-CoT, e.g., “Let’s think step by step.”, is first applied to the test sample before selecting demonstrations to obtain a reasoning path. The training examples that are most semantically similar to the generated reasoning path are then selected as demonstrations. They are prepended to the test sample for inference. Note that IDS ensures that the generated answer is accompanied by the reasoning path through designed prompts. The new reasoning path is then used for extracting another set of demonstration examples by semantic similarity in the next iteration. After a few iterations, IDS adopts majority voting to obtain the final result. Empirical results on tasks spanning mathematical reasoning, commonsense reasoning, logical reasoning, question answering, and topic classification show that IDS can consistently outperform previous ICL demonstration selection baselines. In summary, our main contributions are:
  • We consider both the diversity and similarity dimensions of ICL demonstration selection for LLMs. We identify that the optimal dimension for selecting demonstrations is task-specific and propose Iterative Demonstration Selection (IDS) based on how the test query is answered to fully leverage the merits of both dimensions.
  • With extensive experiments and analysis, we demonstrate the effectiveness of IDS on a variety of tasks.
2 Related Work
This work mainly explores how to select few-shot in-context learning demonstrations for LLMs by leveraging Zero-shot-CoT. In light of this, we review four lines of research that form the basis of this work: few-shot learning, in-context learning basics, demonstration selection for in-context learning, and chain-of-thought reasoning.
2.1 Few-shot Learning
Few-shot learning aims to learn tasks with only a few labeled samples, which results in a big challenge, i.e., over-fitting, for models as they typically require large amounts of data for training. Prior methods to address over-fitting mainly focused on augmenting the few-shot data (Gao et al., 2020; Qin and Joty, 2022), reducing the hypothesis space (Triantafillou et al., 2017; Hu et al., 2018), or optimizing the strategy for searching the best hypothesis (Ravi and Larochelle, 2017; Finn et al., 2017). More recently, LLMs have demonstrated strong few-shot learning ability through in-context learning without any parameter updates (Brown et al., 2020).
2.2 In-context Learning
Brown et al. (2020) first showed that a frozen GPT-3 model can achieve impressive results on a variety of few-shot NLP tasks through conditioning on manually designed prompts consisting of task descriptions and several demonstration examples. Since then many efforts have been made on in-context learning (ICL) (Dong et al., 2022). Chen et al. (2022); Min et al. (2022a); Wei et al. (2023a) demonstrated that the ICL ability of language models can be further improved through self-supervised or supervised training. Some analytical studies attempted to understand what factors affect ICL performance (Zhao et al., 2021; Shin et al., 2022; Wei et al., 2022a; Min et al., 2022b; Yoo et al., 2022; Wei et al., 2023b) and why ICL works (Xie et al., 2022; Olsson et al., 2022; Li et al., 2023b; Pan et al., 2023; Dai et al., 2023). Other ongoing research on ICL has also explored (i) demonstration designing, including demonstration selection (Liu et al., 2022; Rubin et al., 2022; Wang et al., 2023b), demonstration ordering (Lu et al., 2022), and demonstration formatting (Wei et al., 2022b; Wang et al., 2022c; Zhou et al., 2023; Zhang et al., 2023a), (ii) applications of ICL (Ding et al., 2022; Meade et al., 2023; Zheng et al., 2023), and (iii) ICL beyond text (Wang et al., 2023c; Huang et al., 2023; Zhu et al., 2023; Wang et al., 2023a).
2.3 Demonstration Selection for In-context Learning
The performance of ICL has been shown to be highly sensitive to the selection of demonstration examples (Zhao et al., 2021). Existing methods to solve this problem can be mainly divided into two categories. First, unsupervised methods rely on pre-defined metrics. Liu et al. (2022) proposed to select the closest neighbors as demonstrations. In contrast, Levy et al. (2022) selected diverse demonstrations to improve in-context compositional generalization. More recent studies have explored leveraging the output distributions or predictive uncertainty of language models to select few-shot demonstrations (Wu et al., 2022; Nguyen and Wong, 2023; Li and Qiu, 2023; Ma et al., 2023; Ling et al., 2024; Xu and Zhang, 2024) or self-generating demonstrations (Chen et al., 2023). Second, supervised methods involve model training. Rubin et al. (2022); Ye et al. (2023a); Li et al. (2023a); Luo et al. (2023); Wang et al. (2024) proposed to learn to retrieve demonstration examples. Wang et al. (2023b) posited LMs as implicit topic models to facilitate demonstration selection. In addition, some studies (Zhang et al., 2022; Scarlatos and Lan, 2023) attempted to select demonstrations based on reinforcement learning. However, it is still a common practice to randomly select examples or select examples that are semantically similar to the test sample as demonstrations for LLMs for which parameters or detailed output distributions are not available (Sun et al., 2022). Several methods investigated the combination of diversity and similarity in different scenarios, e.g., prompting with explanations (Ye et al., 2023b), choosing examples for annotation (Su et al., 2023) and exploring compositional generalization (An et al., 2023a). Nevertheless, it remains unclear to us how to determine and leverage the optimal dimension for different tasks in ICL and how the reason for answering the test sample benefits the balance between the two dimensions, which motivates us to propose our simple but effective approach (IDS).
2.4 Chain-of-Thought Reasoning
Chain-of-thought (CoT) reasoning induces LLMs to produce intermediate reasoning steps before generating the final answer (Wei et al., 2022b). Depending on whether there are manually designed demonstrations, current CoT reasoning methods mainly include Manual-CoT and Zero-shot-CoT. In Manual-CoT, human-labeled reasoning paths are used to perform CoT reasoning (Wei et al., 2022b; Zhou et al., 2022; Wang et al., 2022b; Li et al., 2022; Wang et al., 2022a). In contrast, LLMs leverage self-generated rationales for reasoning in Zero-shot-CoT (Kojima et al., 2022; Zelikman et al., 2022; Zhang et al., 2023a; Diao et al., 2023). The ongoing research on CoT reasoning has also explored (i) multimodal reasoning (Zhang et al., 2023b; Wu et al., 2023), (ii) distilling knowledge from LLMs (Ho et al., 2022; Fu et al., 2023), and (iii) iterative optimization (Shinn et al., 2023; Madaan et al., 2023; Paul et al., 2023).
3 Problem Formulation
Given the test set $\mathcal{D}_{\text{test}}$ and the training set $\mathcal{D}_{\text{train}}$, the goal of ICL demonstration selection is to find an optimal subset $\mathcal{S}=\{(x_{1},y_{1}),...,(x_{k},y_{k})\}$ ($k$-shot) of $\mathcal{D}_{\text{train}}$ as demonstration examples for each test sample $(\hat{x}_{i},\hat{y}_{i})$ to maximize the overall task performance on $\mathcal{D}_{\text{test}}$. More formally, the optimal selection method $\tilde{h}$ is defined as:
$\tilde{h}=\mathop{\arg\max}\limits_{h\in\mathcal{H}}\sum_{i=1}^{|\mathcal{D}_{\text{test}}|}\delta_{\text{LLM}([h(\mathcal{D}_{\text{train}},\hat{x}_{i},\hat{y}_{i}),\hat{x}_{i}]),\hat{y}_{i}}$ (1)
where $\mathcal{H}$ is the hypothesis space for searching demonstration examples, $h(\mathcal{D}_{\text{train}},\hat{x}_{i},\hat{y}_{i})$ refers to demonstrations selected for $(\hat{x}_{i},\hat{y}_{i})$ using $h$, [,][,][ , ] stands for concatenation, and $\delta_{a,b}$ is the Kronecker delta function: $\delta_{a,b}=1$ if $a$ equals $b$, otherwise $\delta_{a,b}=0$. In this work, we aim to find the optimal method $\tilde{h}$ by leveraging Zero-shot-CoT.
4 What Makes Good In-Context Demonstrations?
CommonsenseQA BoolQ AGNews
Similar-ICL-Consistency (Similarity) 76.0 85.0 90.0
Random-ICL-Voting (Diversity) 79.0 84.0 88.0
Table 1: Results of different methods on CommonsenseQA, BoolQ and AGNews. The optimal dimension for selecting ICL demonstrations is task-specific.
Similar-ICL-Consistency Random-ICL-Voting
Which choice is the correct answer to the question?
Which choice is the correct answer to the question?
Examples: Question: If you have cleaned off dust here it may be difficult to do your homework where? Answer Choices: (A) desktop (B) closet (C) most buildings (D) surface of earth (E) stove Answer: A Question: Where is dust likely to be under? Answer Choices: (A) closet (B) ground (C) windowsill (D) attic (E) carpet Answer: E Question: Where would you find a dustbin that is being used? Answer Choices: (A) utility closet (B) ground (C) cupboard (D) broom closet (E) kitchen Answer: E Question: Dust accumulates where? Answer Choices: (A) ceiling (B) library (C) surface of earth (D) most buildings (E) desktop Answer: D Examples: Question: She had a busy schedule, she had to run errands and pick up the kids the second she did what? Answer Choices: (A) make time for (B) take money (C) go outdoors (D) leave work (E) field Answer: D Question: What is the worst outcome of an injury? Answer Choices: (A) cause death (B) cause bleeding (C) falling down (D) become infected (E) claim insurance Answer: A Question: Mom said that Sarah should stay in bed until she was able to go to school again. What did mom say to Sarah when she tried to get up? Answer Choices: (A) you’re sick (B) were sick (C) more rest (D) rest more (E) get back under the covers Answer: A Question: John got a raise, but he lost rank. Overall, it was a good what? Answer Choices: (A) demotion (B) push down (C) go off strike (D) lower (E) go off strike Answer: A
The response should follow the format: Answer: {A, B, C, D or E} The response should follow the format: Answer: {A, B, C, D or E}
Here is the test data. Here is the test data.
Question: John wanted to clean all of the dust out of his place before settling down to watch his favorite shows. What might he hardest do dust? Answer Choices: (A) closet (B) under the bed (C) television (D) attic (E) most buildings Question: John wanted to clean all of the dust out of his place before settling down to watch his favorite shows. What might he hardest do dust? Answer Choices: (A) closet (B) under the bed (C) television (D) attic (E) most buildings
[…]
Table 2: Examples of Similar-ICL-Consistency (first decoding path) and Random-ICL-Voting (first run) for constructing demonstration examples. The upper part is the input to LLMs, including few-shot demonstrations, and the lower part is the predicted answer. Similar-ICL-Consistency gives the wrong answer ‘most buildings’ which is actually the output of the final demonstration example, indicating that the decision process of the model is misled by this similar sample.
As demonstrated in previous work (Zhao et al., 2021), the overall task performance is highly sensitive to the selection method $h$. Different sets of demonstration examples can yield significantly different performance. For example, Zhang et al. (2022) show that the minimum and maximum ICL performance due to random sampling differs by $>30\%$ on 4 classification tasks, which emphasizes the importance of selecting good demonstrations for LLMs.
A natural question is: what makes good in-context demonstrations? For LLMs, it is still a common practice to select a subset $\mathcal{S}$ consisting of examples that are diverse or semantically similar to the test sample as demonstrations, i.e., considering the diversity or similarity of $\mathcal{S}$. To investigate whether one dimension is consistently better than the other one across different tasks, we conduct some pilot experiments on CommonsenseQA (Talmor et al., 2019), BoolQ (Clark et al., 2019) and AGNews (Zhang et al., 2015). Specifically, we randomly sample $100$ examples from the original test set for experiments and conduct $4$-shot learning using GPT-3.5 (gpt-3.5-turbo).
Following Zhang et al. (2023a), we use Sentence-BERT (Reimers and Gurevych, 2019) to encode all samples. For each test sample, the Similar-ICL method selects the top-$4$ similar training data based on cosine similarity while the Random-ICL method randomly samples $4$ training examples as few-shot demonstrations. Inspired by Wang et al. (2022b), we apply self-consistency with $3$ decoding paths (temperature 0.7) to Similar-ICL (named Similar-ICL-Consistency) and run Random-ICL $3$ times before majority voting (named Random-ICL-Voting) to improve the robustness.
The results of different methods on four datasets are reported in Table 1. We can observe that the diversity dimension outperforms the similarity dimension on CommonsenseQA while the similarity dimension is superior to the diversity dimension on BoolQ and AGNews. Therefore, the optimal dimension for selecting demonstration examples is task-specific. Thus, it is unreasonable to claim that one dimension is consistently better than the other one in ICL demonstration selection.
Intuitively, semantically similar examples can help the model correctly answer the test query as they might share similar input-output patterns with the test sample which could unleash GPT-3.5’s power of text generation. To further understand why the similarity dimension underperforms the diversity dimension on CommonsenseQA, we present a case study in Table 2. We can see that the answer of the final demonstration example extracted by Similar-ICL-Consistency, i.e., ‘most buildings’ is also in the options list of the test sample, which misleads the decision process of the model, leading to a wrong answer. In addition, the selected demonstrations might not include enough important information as high similarity also results in redundancy.
Considering the strengths and weaknesses of both dimensions, we aim to design a method that can select demonstration examples that are diverse (minimizing misleading information) but still strongly correlated with the test sample, which is introduced in the next section.
5 Iterative Demonstration Selection
Figure 2: Illustration of our proposed Iterative Demonstration Selection (IDS). IDS first applies Zero-shot-CoT to the test sample to obtain a reasoning path, which is then used to select few-shot demonstrations from training examples through KNN. The selected demonstration examples are prepended to the test sample for ICL. To obtain the new reasoning path for extracting another set of demonstrations in the next iteration, an instruction for output format is inserted before the test sample. After several iterations, IDS uses majority voting to obtain the final result.
Based on the observations and considerations in Section 4, we introduce Iterative Demonstration Selection (IDS) for ICL demonstration selection by leveraging how the test sample is answered (see Fig. 2 for an illustration). Intuitively, the demonstrations that are similar to the reason for answering a sample are strongly correlated with this sample. Therefore, we propose to incorporate zero-shot chain-of-thought reasoning (Zero-shot-CoT) into IDS to iteratively select demonstration examples that are diverse but still have a strong correlation with the test sample.
Specifically, for each test sample $\hat{x}_{i}$, IDS mainly consists of four steps:
  1. 1.
    We apply Zero-shot-CoT, i.e., “Let’s think step by step.” to the test sample $\hat{x}_{i}$ before selecting demonstrations to obtain a reasoning path $\mathrm{R}$.
  2. 2.
    The reasoning path $\mathrm{R}$ is then used to select top-$k$ ($k$ is the number of shot) most semantically similar training examples $\{(x_{1},y_{1}),...,(x_{k},y_{k})\}$ as few-shot demonstrations. We use Sentence-BERT (Reimers and Gurevych, 2019) to encode the reasoning path $\mathrm{R}$ and training examples to obtain the contextual representations and use cosine similarity to measure the similarity between representations.
  3. 3.
[…]
  4. 4.
    After $q$ rounds of iterations between Step 2 and 3, we adopt majority voting on all $\hat{\mathrm{A}}$ to obtain the final result $\hat{\mathrm{A}}_{final}$.
Algorithm 1 Selection process of IDS
1:Training set $\mathcal{D}_{\text{train}}$, test set $\mathcal{D}_{\text{test}}$, $\text{LLM}_{\theta}$, number of demonstrations $k$, number of iterations $q$ and answer set $\hat{\mathrm{A}}_{all}=\emptyset$
2:Encode all samples in $\mathcal{D}_{\text{train}}$ using Sentence-BERT $\triangleright$ Encode training set
3:for $\hat{x}_{i}$ in $\mathcal{D}_{\text{test}}$ do
4:     Apply Zero-shot-CoT to $\hat{x}_{i}$ to obtain the reasoning path $\mathrm{R}$triangleright Zero-shot-CoT
5:     for $j=1,\ldots,q$ do
6:         Encode $\mathrm{R}$ using Sentence-BERT $\triangleright$ Encode reasoning path
7:         Use $\mathrm{R}$ to select top-$k$ most similar examples $\mathcal{S}=\{(x_{1},y_{1}),...,(x_{k},y_{k})\}$ from $\mathcal{D}_{\text{train}}$ as demonstrations $\triangleright$ KNN selection
8:         ($\hat{\mathrm{A}}$, $\hat{\mathrm{R}}$) = $\text{LLM}_{\theta}(S,\hat{x}_{i})$triangleright ICL with Zero-shot-CoT
9:         $\mathrm{R}=\hat{\mathrm{R}}$, $\hat{\mathrm{A}}_{all}=\hat{\mathrm{A}}_{all}\cup\{\hat{\mathrm{A}}\}$triangleright Update reasoning path and answer set
10:     end for
11:     Adopt majority voting for $\hat{\mathrm{A}}_{all}$ to obtain the final result $\hat{\mathrm{A}}_{final}$ for the test sample $\hat{x}_{i}$triangleright Majority voting
12:end for
Obviously, the selected demonstration examples are strongly correlated with the original test sample, i.e., achieving similarity, as they are selected by the generated reasoning paths (see Section A.4 for quantitative analysis of reasoning paths). And they can be different during iterations to achieve diversity because the reasoning paths vary in different iterations. Note that there is no reasoning path in few-shot demonstrations (as shown in the green part in Fig. 2). The reasoning path only exists in the output of LLMs.
In addition, we illustrate the whole selection process in Alg. 1 and show the instructions and input formats of different types of tasks for ICL in Section A.1.
Method BoolQ GSM8K MATH CommonsenseQA LogiQA AGNews Average
Vote-$k$ $86.7_{\pm 0.7}$ $76.5_{\pm 0.5}$ $35.7_{\pm 0.2}$ $75.2_{\pm 0.3}$ $45.4_{\pm 0.3}$ $88.1_{\pm 1.2}$ $67.9_{\pm 0.2}$
MMR $86.4_{\pm 0.8}$ $75.5_{\pm 0.7}$ $34.8_{\pm 0.3}$ $74.9_{\pm 0.2}$ $44.7_{\pm 0.3}$ $87.6_{\pm 1.1}$ $67.3_{\pm 0.3}$
G-fair-Prompting $84.8_{\pm 0.7}$ $76.9_{\pm 0.6}$ $34.6_{\pm 0.3}$ $75.5_{\pm 0.3}$ $43.8_{\pm 0.4}$ $88.9_{\pm 1.0}$ $67.4_{\pm 0.2}$
Skill-KNN $85.9_{\pm 0.5}$ $76.5_{\pm 0.3}$ $35.1_{\pm 0.2}$ $75.2_{\pm 0.2}$ $44.6_{\pm 0.2}$ $88.7_{\pm 0.9}$ $67.7_{\pm 0.1}$
Top-$k$-Consistency $87.1_{\pm 0.2}$ $76.1_{\pm 0.5}$ $35.6_{\pm 0.3}$ $74.5_{\pm 0.2}$ $45.7_{\pm 0.4}$ $89.3_{\pm 0.8}$ $68.1_{\pm 0.1}$
Random-Voting $87.3_{\pm 0.6}$ $75.6_{\pm 0.4}$ $35.4_{\pm 0.1}$ $77.0_{\pm 0.2}$ $45.1_{\pm 0.3}$ $87.0_{\pm 1.6}$ $67.9_{\pm 0.2}$
Cluster-Voting $86.4_{\pm 0.7}$ $76.8_{\pm 0.3}$ $34.9_{\pm 0.4}$ $76.5_{\pm 0.3}$ $44.1_{\pm 0.3}$ $86.8_{\pm 1.2}$ $67.6_{\pm 0.3}$
IDS $\mathbf{87.8}_{\pm 0.8}$ $\mathbf{78.5}_{\pm 0.4}$ $\mathbf{37.5}_{\pm 0.2}$ $\mathbf{78.1}_{\pm 0.1}$ $\mathbf{46.9}_{\pm 0.2}$ $\mathbf{89.8}_{\pm 0.8}$ $\mathbf{69.8}_{\pm 0.1}$
[…]
6 Experiments
In this section, we first describe the tasks and datasets, and then introduce methods compared in our work. Finally, we present the experimental results.
6.1 Experimental Setup
Tasks and Datasets
We mainly investigate 6 different datasets covering 5 representative task categories: mathematical reasoning (GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021)), commonsense reasoning (CommonsenseQA (Talmor et al., 2019)), logical reasoning (LogiQA (Liu et al., 2020)), question answering (BoolQ (Clark et al., 2019)) and topic classification (AGNews (Zhang et al., 2015)). For each dataset, we randomly sample at most 10000 examples from the original training set as $\mathcal{D}_{\text{train}}$ and at most 2000 test examples as $\mathcal{D}_{\text{test}}$ for evaluating the performance of selected demonstrations. The detailed information of different datasets is shown in Section A.2. To reduce the randomness, we run every experiment five times with different random seeds (resulting in different training and test samples if not using the whole set) and report the average results. Without specification, we use $k$ = 4 number of demonstrations following Wang et al. (2023b) and set the number of iterations $q$ to $3$.
Methods Compared
We mainly use GPT-3.5 (gpt-3.5-turbo) as the LLM and compare our IDS with the following methods in the experiments for selecting ICL demonstrations:
  • Top-$k$-Consistency (Liu et al., 2022) selects the top-$k$ semantically similar examples from the training set $\mathcal{D}_{\text{train}}$ as demonstrations for each test sample and applies self-consistency (Wang et al., 2022b) with $q$ decoding paths (temperature 0.7) to match the number of iterations. Following Zhang et al. (2023a), all samples are encoded by Sentence-BERT (Reimers and Gurevych, 2019) to obtain contextual representations for calculating the cosine similarity.
  • Random-Voting randomly selects $k$ examples from $\mathcal{D}_{\text{train}}$ as few-shot demonstrations for every test sample and runs experiments $q$ times before majority voting.
  • Cluster-Voting partitions $\mathcal{D}_{\text{train}}$ into $k$ clusters and selects a representative example from each cluster to form demonstrations. Following Zhang et al. (2023a), we choose the sample closest to the centroid in each cluster as the representative example. Same as Random-Voting, after running experiments $q$ times, Cluster-Voting adopts majority voting to obtain the final result.
[…]
Top-$k$-Consistency IDS Random-Voting
Average Similarity Score 0.68 0.48 0.32
Table 4: Average similarity scores between test examples and the corresponding selected demonstrations of three methods (Top-$k$-Consistency, IDS and Random-Voting).
[…]
6.2 Main Results
Table 3 shows the average performance scores of different methods on all investigated datasets. From the results, we can observe that
[…]
To delve deeper into how different dimensions are leveraged in selected demonstrations, we report the average similarity scores between test samples and the corresponding demonstrations of different methods in Table 4. Specifically, we randomly select 500 test examples for each dataset and use Sentence-BERT to obtain contextual representations for calculating similarity scores. We can see that the average similarity score of IDS is between that of Top-$k$-Consistency and Random-Voting, indicating that it can indeed strike a balance between two selection dimensions (see Section A.5 for more analysis on the diversity of the selected demonstration examples).
6.3 Analysis
Different Numbers of Iterations
Our experiments and analysis so far use $q=3$ iterations. To verify whether the performance gain of IDS is consistent across different numbers of iterations, we conduct controlled experiments with $q=\{1,5,7\}$. The average results of the 6 datasets with a randomly selected seed are reported in Fig. 3. IDS consistently outperforms the best baseline Top-$k$-Consistency with different $q$ (even $q=1$, i.e., without voting), emphasizing the importance of rationales in selecting demonstration examples. Interestingly, the performance of ICL does not always improve with the number of iterations, which might be because increased iterations can also lead to unnecessary noise; we provide an in-depth analysis in Section A.6.
GPT-3.5 GPT-4
Top-$k$-Consistency 68.3 73.9
IDS 69.9 75.4
[…]
Robustness to Model Types
To demonstrate the robustness of IDS to model types, we conduct controlled experiments with GPT-4. Specifically, we randomly select one seed and sample $200$ test examples per dataset for experiments due to the expensive cost. From the average results shown in Table 5, we can observe that IDS still achieves better performance than Top-$k$-Consistency when using GPT-4 as the LLM, showing its robustness to different LLMs.
Generalization to Open-source LLMs
To better verify the generalization ability of IDS, we use vLLM (Kwon et al., 2023) to serve Llama-2-chat models (Touvron et al., 2023) for experiments and compare IDS with Top-$k$-Consistency on two datasets: BoolQ and GSM8K. We randomly sample 500 test examples for experiments and report the results in Table 6, which demonstrates that IDS can successfully generalize to open-source LLMs of different sizes.
BoolQ GSM8K
7B 13B 70B 7B 13B 70B
Top-$k$-Consistency 77.1 81.3 84.2 14.6 24.8 49.6
IDS 78.5 82.2 85.4 16.6 27.1 51.4
[…]
Figure 4: Several case studies of model responses. We color correct outputs in green, and wrong outputs in red.
Case Study
To further understand the advantage of IDS, we show several cases in Fig. 4. As shown in the upper part of the figure, IDS can iteratively select more diverse demonstration examples than Top-$k$-Consistency which may be able to correct errors from previous iterations. Compared with Random-Voting, IDS can find examples that share more similar input-output patterns with the test sample to induce the LLM to generate correct answers (the lower part of the figure).
In addition, we show the results with different numbers of demonstrations, the robustness of IDS to different embedding models and Zero-shot-CoT triggers, and the results on two additional datasets in Appendix A.7 $\sim$ A.10, respectively.
7 Conclusion
In this work, we have introduced Iterative Demonstration Selection (IDS) that can iteratively select examples that are diverse but still strongly correlate with the test sample as demonstrations to improve the performance of in-context learning (ICL) by leveraging the rationale for answering the test sample. Extensive experimental results and analysis show that IDS can consistently outperform previous ICL demonstration selection baselines.
Limitations
This work has several limitations. First, due to the inference cost of ChatGPT, we do not conduct experiments on the entire test set. Besides, we include 6 datasets covering 5 different task types in this work. A further improvement could be to explore more diverse types of tasks.
Appendix A Appendix
A.1 Instructions and Input Formats of Different Tasks
We show the instructions and input formats of different types of tasks for in-context learning in Fig. 5.
Figure 5: Instructions and input formats of five different categories of tasks (topic classification, question answering, commonsense reasoning, logical reasoning, and mathematical reasoning) for ICL. For Zero-shot-CoT in the first step of IDS, there is no demonstration example and the instruction “Here is the test data.”.
A.2 Datasets Information
BoolQ GSM8K MATH CommonsenseQA LogiQA AGNews
# Training Samples 9427 (full) 7473(full) 5000 9741 (full) 7376(full) 10000
# Test Samples 2000 1000 1000 1221 (full) 500 1000
Table 7: Deailed information of different datasets. # refers to ‘the number of’ and ‘full’ means the whole set. Note that different random seeds do not result in different samples if the whole set is used.
We show the detailed information of different datasets in Table 7.
A.3 Details of Baselines
In this work, we compare IDS with the following latest ICL demonstration selection approaches:
  • Vote-$k$ (Su et al., 2023) is an unsupervised, graph-based selective annotation method used for selecting and annotating diverse, representative examples. The annotated examples then serve as a pool for demonstration retrieval.
  • MMR (Ye et al., 2023b) proposes a maximal marginal relevance-based approach for demonstration selection.
  • G-fair-Prompting (Ma et al., 2023) leverages greedy search to select the example with the highest fairness score at each step.
  • Skill-KNN (An et al., 2023b) generates skill-based descriptions for test queries and then uses these descriptions to select similar examples as demonstrations.
A.4 Measure of Reasoning Path Correlation
$\text{score}_{\text{{reason}}}$ $\text{score}_{\text{{random}}}$ $\text{score}_{\text{{similar}}}$
Average Similarity Score 0.59 0.32 0.68
Table 8: Comparison between different average similarity scores.
We report the average similarity score between test samples and the corresponding generated reasoning paths ($\text{score}_{\text{{reason}}}$), the average similarity score between test samples and randomly selected training examples ($\text{score}_{\text{{random}}}$), and the average similarity score between test samples and the most similar training examples ($\text{score}_{\text{{similar}}}$) in Table 8. For each dataset, we randomly select 500 test samples and use Sentence-BERT for similarity calculation. We can observe that $\text{score}_{\text{{reason}}}$ is slightly worse than $\text{score}_{\text{{similar}}}$ and much higher than $\text{score}_{\text{{random}}}$, indicating that the generated reasoning path is indeed strongly correlated with the test sample.
A.5 Analysis on Demonstration Diversity
Top-$k$-Consistency IDS
Average Pairwise Similarity 0.55 0.39
Table 9: Comparison of average pairwise similarity scores of demonstrations selected by different methods.
In addition to the average similarity score between test samples and demonstrations, we further calculate the following metrics for IDS and Top-$k$-Consistency:
$Q_{S}=\sum_{1\leq i<j\leq|S|}g(S_{i},S_{j})/C(|S|,2)$ (2)
where $S$ is the set of the selected demonstration examples, and $g$ is the function of measuring similarity. $Q$ calculates the average pairwise similarity score of the demonstrations, which can be used to reflect whether they are diverse from each other. As can be seen from the results in Table 9, the average pairwise similarity score of IDS is much lower than that of Top-$k$-Consistency, verifying the diversity of demonstration examples selected by IDS.
A.6 Noise Caused by Increased Iterations
Iteration 5 7
[…]
Table 10: Comparison between different iterations.
As observed from Fig. 3, the performance of ICL does not always improve with the number of iterations. We speculate that this is because too many iterations may also lead to unnecessary noise. As the number of iterations increases, the demonstrations selected in the latest iteration are more likely to have been chosen in previous iterations. Therefore, if these demonstrations result in wrong answers in previous iterations, these errors may be propagated to later iterations, i.e., unnecessary noise caused by increased iterations. To better verify our hypothesis, we calculate (i) the proportion of demonstrations selected in iteration 5 or 7 that were also chosen in previous iterations ($\text{Prop}_{\text{pre}}$), and (ii) the proportion of demonstrations selected in iteration 5 or 7 that were chosen in previous iterations and resulted in wrong answers ($\text{Prop}_{\text{pre}}^{\text{wrong}}$). We can see from Table 10 that the results of the 7th iteration are much higher than those of the 5th iteration, indicating the correctness of our claim.
A.7 Different Numbers of Demonstrations
2 4 6 8
Top-$k$-Consistency 68.0 68.3 68.5 68.4
IDS 69.4 69.9 69.9 69.7
[…]
While we use $k=4$ demonstration examples for all experiments, we also evaluate the effectiveness of IDS with different $k$. We randomly choose one seed for experiments and report the average results of the 6 datasets in Table 11. We can see that IDS consistently outperforms Top-$k$-Consistency with different numbers of demonstrations. In addition, more demonstrations do not guarantee better ICL performance, which is consistent with the observation in Wang et al. (2023b).
BoolQ CommonsenseQA GSM8K
Top-$k$-Consistency 86.0 75.4 75.8
IDS 87.2 78.0 77.6
[…]
A.8 Robustness to Embedding Models
Instead of using Sentence-BERT, we also explore adopting the OpenAI embedding model (text-embedding-ada-002) as the encoder. Specifically, we conduct experiments on 3 datasets: BoolQ, CommonsenseQA and GSM8K. For each dataset, we randomly sample 500 test examples and compare IDS with the baseline Top-$k$-Consistency. The results reported in Table 12 demonstrate IDS’s robustness to different embedding models.
Default Trigger1 Trigger2
IDS 70.1 70.3 70.0
[…]
A.9 Robustness to Zero-shot-CoT Triggers
To verify the robustness of IDS to Zero-shot-CoT triggers, we conduct controlled experiments with two new triggers: “Let’s work this out in a step by step way to be sure we have the right answer.” (Trigger1) and “Let’s solve this problem step by step” (Trigger2). Specifically, we randomly sample 500 test examples per dataset for experiments and report the average results in Table 13, which demonstrates that IDS is indeed robust to different Zero-shot-CoT triggers.
A.10 Two Additional Datasets
MNLI Emotion
Top-$k$-Consistency 65.7 58.1
IDS 67.4 60.3
Table 14: Results on two additional datasets.
To better demonstrate the generalization ability of IDS, we further conduct experiments on two additional datasets: MNLI (natural language inference) (Williams et al., 2018) and Emotion (emotion classification) (Saravia et al., 2018). The comparison between IDS and the baseline Top-$k$-Consistency is shown in Table 14, which verifies the strong generalizability of IDS.

### v4

@@
 Despite the effectiveness, the performance of ICL has been shown to be highly sensitive to the selection of demonstration examples (Zhao et al., 2021).
-Different sets of demonstrations can yield performance ranging from nearly random to comparable with state-of-the-art models (Gao et al., 2021; Lu et al., 2022).
+Different sets of demonstrations can yield performance ranging from nearly random to comparable with state-of-the-art models (Gao et al., 2021; Lu et al., 2022; Qin et al., 2023b).
 To alleviate the above issue, researchers in ICL have proposed a number of methods to select a set of examples as few-shot demonstrations (Rubin et al., 2022; Liu et al., 2022; Li and Qiu, 2023; Wang et al., 2023b; Li et al., 2023a; Ma et al., 2023; An et al., 2023b).
@@
 Few-shot learning aims to learn tasks with only a few labeled samples, which results in a big challenge, i.e., over-fitting, for models as they typically require large amounts of data for training.
-Prior methods to address over-fitting mainly focused on augmenting the few-shot data (Gao et al., 2020; Qin and Joty, 2022), reducing the hypothesis space (Triantafillou et al., 2017; Hu et al., 2018), or optimizing the strategy for searching the best hypothesis (Ravi and Larochelle, 2017; Finn et al., 2017).
+Prior methods to address over-fitting mainly focused on augmenting the few-shot data (Gao et al., 2020; Qin and Joty, 2022a, b), reducing the hypothesis space (Triantafillou et al., 2017; Hu et al., 2018; Qin et al., 2023a), or optimizing the strategy for searching the best hypothesis (Ravi and Larochelle, 2017; Finn et al., 2017; Xia et al., 2024).
 More recently, LLMs have demonstrated strong few-shot learning ability through in-context learning without any parameter updates (Brown et al., 2020).
@@
 Depending on whether there are manually designed demonstrations, current CoT reasoning methods mainly include Manual-CoT and Zero-shot-CoT.
-In Manual-CoT, human-labeled reasoning paths are used to perform CoT reasoning (Wei et al., 2022b; Zhou et al., 2022; Wang et al., 2022b; Li et al., 2022; Wang et al., 2022a).
-In contrast, LLMs leverage self-generated rationales for reasoning in Zero-shot-CoT (Kojima et al., 2022; Zelikman et al., 2022; Zhang et al., 2023a; Diao et al., 2023).
-The ongoing research on CoT reasoning has also explored (i) multimodal reasoning (Zhang et al., 2023b; Wu et al., 2023), (ii) distilling knowledge from LLMs (Ho et al., 2022; Fu et al., 2023), and (iii) iterative optimization (Shinn et al., 2023; Madaan et al., 2023; Paul et al., 2023).
+In Manual-CoT, human-labeled reasoning paths are used to perform CoT reasoning (Wei et al., 2022b; Zhou et al., 2022; Wang et al., 2022b; Li et al., 2022; Wang et al., 2022a; Zhao et al., 2023).
+In contrast, LLMs leverage self-generated rationales for reasoning in Zero-shot-CoT (Kojima et al., 2022; Zelikman et al., 2022; Zhang et al., 2023a; Diao et al., 2023; Qin et al., 2024).
+The ongoing research on CoT reasoning has also explored (i) multimodal reasoning (Zhang et al., 2023b; Wu et al., 2023), (ii) distilling knowledge from LLMs (Ho et al., 2022; Fu et al., 2023), and (iii) iterative optimization (Shinn et al., 2023; Madaan et al., 2023; Paul et al., 2023; Jiao et al., 2024).
 3 Problem Formulation
@@
 Compared with Random-Voting, IDS can find examples that share more similar input-output patterns with the test sample to induce the LLM to generate correct answers (the lower part of the figure).
-In addition, we show the results with different numbers of demonstrations, the robustness of IDS to different embedding models and Zero-shot-CoT triggers, and the results on two additional datasets in Appendix A.7 $\sim$ A.10, respectively.
+In addition, we show the results with different numbers of demonstrations, the robustness of IDS to different embedding models and Zero-shot-CoT triggers in Appendix A.7 $\sim$ A.9, respectively.
 7 Conclusion
@@
 Specifically, we randomly sample 500 test examples per dataset for experiments and report the average results in Table 13, which demonstrates that IDS is indeed robust to different Zero-shot-CoT triggers.
-A.10 Two Additional Datasets
-MNLI Emotion
-Top-$k$-Consistency 65.7 58.1
-IDS 67.4 60.3
-Table 14: Results on two additional datasets.
-To better demonstrate the generalization ability of IDS, we further conduct experiments on two additional datasets: MNLI (natural language inference) (Williams et al., 2018) and Emotion (emotion classification) (Saravia et al., 2018).
-The comparison between IDS and the baseline Top-$k$-Consistency is shown in Table 14, which verifies the strong generalizability of IDS.